# Differentially Private 2D Human Pose Estimation

### v2

@@ 4.2.3. Improved Results via Projection
$\epsilon=1.0$ 4.40 19.75 8.40 10.59 22.54 10.92 6.90 13.48 0.59
$\epsilon=3.0$ 15.45 21.79 8.23 10.57 23.21 12.59 4.65 14.89 0.66
$\epsilon=5.0$ 15.93 14.83 6.34 9.90 23.11 12.33 7.25 13.16 0.59
$\epsilon=7.0$ 0.75 17.02 14.79 9.39 23.58 15.76 4.91 14.09 0.55
Figure 1 compares the mean accuracy improvements achieved by introducing projection to standard DP-SGD across broad range of privacy budgets and clipping thresholds.
Figure 1. Comparison of PCKh@0.5 with and without projection across different training strategies with varied epsilons and clipping norm.
The integration of gradient projection in DP-SGD consistently boosts performance across nearly all experimental configurations, demonstrating its effectiveness in enhancing differential private 2D-HPE training. These improvements are most substantial at tight privacy budgets (lower $\epsilon$ values), where projection significantly mitigates the inherent accuracy-privacy trade-off typically observed in DP training. We observe a particularly strong impact of projection with higher clipping thresholds (e.g., $C=1.0$ and $C=0.1$). Here, projection dramatically boost performance under stringent privacy constraints. For example, at $\epsilon=0.4$ and $C=1.0$ accuracy increases more than 25% (from 8.36% to 36.99%), highlighting the significant advantage gained by incorporating projection under challenging privacy settings.
The benefits of projection become less pronounced at the lowest clipping threshold ($C=0.01$) and more relaxed privacy budgets (higher $\epsilon$ values). In these cases, while projection still achieves improved accuracy, the gains plateau, suggesting that extremely tight gradient clipping already constraints gradients efficiently in low privacy settings. Interestingly, projection proves especially beneficial when fine-tuning from scratch with $C=0.01$, substantially narrowing the performance gap typically observed compared to fine-tuning from pretrained models. Conversely, projection’s benefit are observed only at low $espilon$ values when training entirely from scratch with $C=0.01$. These results might indicates how challenging is to train DP models from scratch.
In summary, our projection-enhanced DP-SGD approach for 2D-HPE consistently improves accuracy under DP constraints, with the greatest benefits emerging under stricter privacy conditions and moderate gradient clipping thresholds. These findings provide valuable guidance for applying DP in HPE and other sensitive computer vision, deep learning tasks.
5. Conclusion
Our work presents the first differentially private (DP) approach to 2D human pose estimation (HPE), addressing critical privacy concerns while maintaining utility for multimedia applications. By integrating Projected DP-SGD (PDP-SGD) and adapting TinyViT with a coordinate classification head, we achieved remarkable accuracy (78.48% PCKh@0.5 at $epsilon=0.2$) even under strict privacy constraints, significantly outperforming standard DP-SGD procedure. Our comprehensive experiments on the MPII dataset revealed that fine-tuning with frozen initial layers and a low clipping thresholds (C=0.01) provides the optimal balance between privacy and performance, with notable improvements for challenging joints. Unlike traditional anonymization techniques that compromise data utility, our method maintains analytical value while providing formal privacy guarantees, laying a solid foundation for privacy-preserving HPE in sensitive domains such as healthcare and activity recognition. This work demonstrates that DP pose estimation is not only feasible but can approach non-private performance levels while ensuring strong privacy protection.
6. Acknowledgements
The authors acknowledge funding from EPSRC (EP/W01212X/1), Royal Society (RGS/R2/ 212199) and Academy of Medical Sciences (NGR1/1678).

### v3

@@ Appendix A Implementation Details
Table 4: MPII Results: Feature DP-SGD.
Finetuning
C = 0.01
$\epsilon=0.2$ 87.04 89.28 75.78 66.54 75.89 65.04 58.48 75.47 10.64
$\epsilon=0.4$ 90.76 91.30 78.51 69.95 78.55 68.85 63.86 78.60 12.71
$\epsilon=0.6$ 91.47 91.95 79.63 71.99 79.87 71.00 65.45 79.90 13.62
$\epsilon=0.8$ 92.16 92.15 79.85 72.81 80.41 71.55 66.30 80.41 14.22
C = 0.1
$\epsilon=0.2$ 52.42 65.78 57.32 27.43 53.31 36.05 26.03 48.99 3.53
$\epsilon=0.4$ 78.04 81.62 64.53 40.43 67.70 45.14 39.58 61.73 5.44
$\epsilon=0.6$ 78.48 84.80 68.48 45.60 69.88 49.89 45.11 65.32 6.47
$\epsilon=0.8$ 82.74 85.21 70.29 48.50 71.39 54.50 47.85 67.60 7.03
C = 1.0
$\epsilon=0.2$ 1.60 9.90 7.86 11.57 14.25 3.00 7.91 9.32 0.33
$\epsilon=0.4$ 11.49 17.05 8.30 11.08 23.89 8.20 9.57 15.01 0.59
$\epsilon=0.6$ 14.84 25.00 12.36 13.62 27.07 10.84 11.29 18.37 0.81
$\epsilon=0.8$ 21.69 29.28 16.16 14.56 26.55 15.66 15.45 22.05 1.01
Finetuning from scratch
C = 0.01
$\epsilon=0.2$ 71.15 57.17 38.71 22.32 47.24 28.39 29.12 43.89 3.48
$\epsilon=0.4$ 78.75 69.40 47.15 33.34 54.86 38.87 37.72 53.00 5.10
$\epsilon=0.6$ 82.44 72.69 52.00 38.19 59.22 41.69 41.14 56.78 6.07
$\epsilon=0.8$ 83.80 74.81 54.88 41.29 60.93 44.35 42.80 58.98 6.70
C = 0.1
$\epsilon=0.2$ 33.94 30.42 19.36 12.83 27.19 15.03 10.82 23.15 1.14
$\epsilon=0.4$ 40.86 37.75 22.57 15.08 31.71 18.09 15.94 28.46 1.53
$\epsilon=0.6$ 47.68 43.27 25.09 16.31 35.35 19.56 17.78 31.90 1.92
$\epsilon=0.8$ 52.25 45.31 27.27 16.57 35.90 19.28 18.47 33.27 2.21
C = 1.0
$\epsilon=0.2$ 12.14 8.85 8.76 10.54 11.72 9.47 4.04 10.03 0.39
$\epsilon=0.4$ 10.06 13.33 10.74 9.92 20.84 12.17 5.48 12.57 0.51
$\epsilon=0.6$ 7.03 12.65 17.11 11.29 21.67 15.35 3.19 12.84 0.52
$\epsilon=0.8$ 13.47 19.74 15.95 9.58 22.95 12.65 8.90 15.69 0.60
Training from scratch
C = 0.01
$\epsilon=0.2$ 6.92 18.29 17.40 10.96 22.62 14.89 3.71 15.17 0.60
$\epsilon=0.4$ 10.57 22.61 18.41 12.15 23.09 14.33 6.83 17.14 0.71
$\epsilon=0.6$ 11.39 22.18 17.16 11.91 24.23 15.90 8.86 17.57 0.82
$\epsilon=0.8$ 14.09 21.31 19.52 11.31 24.32 16.44 7.35 17.74 0.80
C = 0.1
$\epsilon=0.2$ 4.23 0.56 7.82 8.64 19.44 1.91 1.96 7.13 0.26
$\epsilon=0.4$ 0.48 12.62 10.69 8.12 21.15 2.36 0.33 8.87 0.37
$\epsilon=0.6$ 0.48 15.08 10.64 9.00 19.72 4.33 1.20 9.60 0.41
$\epsilon=0.8$ 0.61 13.09 12.94 10.74 22.73 4.94 1.37 11.22 0.48
C = 1.0
$\epsilon=0.2$ 0.31 0.32 0.14 0.00 0.03 0.06 0.12 0.35 0.02
$\epsilon=0.4$ 1.84 3.63 0.02 0.43 0.00 0.00 0.09 0.78 0.02
$\epsilon=0.6$ 0.00 2.62 0.05 4.61 8.34 0.12 0.00 2.39 0.10
$\epsilon=0.8$ 0.14 0.02 0.07 0.00 15.22 0.02 0.26 2.57 0.12
Table 5: MPII Results: Feature Projection DP-SGD.
Finetuning
C = 0.01
$\epsilon=0.2$ 94.17 92.70 82.17 73.29 80.79 76.45 72.60 82.50 19.65
$\epsilon=0.4$ 94.13 92.70 81.46 73.10 80.51 75.44 71.00 82.01 19.23
$\epsilon=0.6$ 94.10 92.70 81.47 72.79 80.87 75.32 71.19 82.01 19.48
$\epsilon=0.8$ 94.27 92.63 81.58 72.50 80.37 75.56 70.88 81.91 19.22
C = 0.1
$\epsilon=0.2$ 93.79 91.85 79.07 71.56 77.43 73.32 68.92 80.24 15.20
$\epsilon=0.4$ 94.03 92.56 80.72 73.19 79.85 74.95 69.96 81.60 17.03
$\epsilon=0.6$ 93.86 92.82 81.17 74.70 79.78 75.68 70.78 82.09 17.74
$\epsilon=0.8$ 94.78 93.21 82.31 74.51 80.63 76.04 71.30 82.62 18.64
C = 1.0
$\epsilon=0.2$ 73.36 65.88 53.72 40.79 56.14 43.44 34.10 54.75 4.05
$\epsilon=0.4$ 86.53 82.17 63.52 54.50 57.68 49.20 43.01 64.02 6.09
$\epsilon=0.6$ 86.66 86.06 66.70 51.94 67.47 51.40 47.02 67.02 6.73
$\epsilon=0.8$ 91.13 89.06 73.91 63.39 62.92 59.60 51.75 71.66 9.37
Finetuning from scratch
C = 0.01
$\epsilon=0.2$ 92.91 88.94 76.55 67.33 77.81 71.11 65.99 78.11 16.82
$\epsilon=0.4$ 92.74 88.55 75.80 65.63 77.48 70.02 65.37 77.41 16.23
$\epsilon=0.6$ 93.21 88.62 75.61 65.15 77.79 69.33 64.52 77.23 16.37
$\epsilon=0.8$ 92.29 87.11 74.11 63.13 76.53 67.46 62.78 75.74 15.83
C = 0.1
$\epsilon=0.2$ 91.95 85.61 71.11 59.78 73.84 64.58 61.36 73.51 14.33
$\epsilon=0.4$ 93.76 88.65 76.97 67.07 77.27 70.30 66.23 78.01 16.93
$\epsilon=0.6$ 93.49 89.08 77.09 67.81 79.09 72.23 67.48 78.86 16.93
$\epsilon=0.8$ 94.03 90.46 78.56 68.58 80.68 72.54 69.08 79.95 17.83
C = 1.0
$\epsilon=0.2$ 4.20 8.02 13.07 9.20 22.99 7.64 7.02 10.98 0.46
$\epsilon=0.4$ 10.54 17.53 13.19 11.50 20.60 7.41 5.12 13.49 0.53
$\epsilon=0.6$ 15.52 20.67 14.91 12.11 21.64 14.99 8.10 16.29 0.78
$\epsilon=0.8$ 20.36 20.92 12.63 10.95 24.98 10.07 8.46 16.29 0.73
Training from scratch
C = 0.01
$\epsilon=0.2$ 16.75 19.53 17.62 12.41 24.84 14.35 9.99 17.34 0.71
$\epsilon=0.4$ 62.65 52.11 34.07 19.03 41.09 27.30 23.74 38.90 3.43
$\epsilon=0.6$ 71.66 60.36 38.86 22.51 46.65 31.51 26.26 44.28 4.39
$\epsilon=0.8$ 67.84 58.95 38.06 21.86 44.95 31.99 27.66 43.39 4.12
C = 0.1
$\epsilon=0.2$ 14.02 17.56 14.69 10.86 20.65 15.84 7.96 15.50 0.72
$\epsilon=0.4$ 11.02 19.51 16.62 11.02 23.61 14.16 10.16 16.42 0.65
$\epsilon=0.6$ 18.21 24.56 19.24 12.34 26.62 15.68 12.49 19.78 0.93
$\epsilon=0.8$ 53.27 46.01 29.37 17.27 35.47 21.20 17.17 33.49 2.33
C = 1.0
$\epsilon=0.2$ 1.13 14.74 9.56 8.19 22.14 14.47 3.54 12.39 0.54
$\epsilon=0.4$ 17.77 14.81 15.85 9.46 22.16 14.83 4.06 15.23 0.70
$\epsilon=0.6$ 18.49 21.08 15.39 11.12 22.54 5.14 5.48 14.67 0.66
$\epsilon=0.8$ 5.83 14.96 14.88 10.66 22.21 13.04 3.83 12.97 0.55
Table 6: HumanART Results: DP-SGD.
Privacy Parameter($\epsilon$) Head Shoulder Elbow Wrist Hip Knee Ankle Mean Mean@0.1 inkoti
Finetuning
C = 0.01
$\epsilon=0.2$ 29.1 67.6 20.0 14.0 30.9 34.0 71.6 28.6 22.0 35.5
$\epsilon=0.4$ 40.7 74.7 39.3 23.8 42.7 45.8 77.1 46.7 32.9 47.9
$\epsilon=0.6$ 37.5 74.7 33.6 20.9 39.4 42.1 77.6 41.4 29.1 43.8
$\epsilon=0.8$ 39.0 75.9 36.0 22.2 40.9 43.5 78.3 43.6 30.1 45.3
C = 0.1
$\epsilon=0.2$ 3.5 19.0 0.0 0.9 4.0 8.7 35.6 0.8 5.0 9.2
$\epsilon=0.4$ 7.7 32.4 0.5 2.3 8.5 14.6 47.7 4.0 8.7 15.3
$\epsilon=0.6$ 10.4 39.3 1.5 3.4 11.3 17.1 51.9 6.2 10.5 17.9
$\epsilon=0.8$ 12.0 43.3 2.4 3.9 13.0 18.4 54.1 7.3 11.4 19.2
Finetuning from scratch
C = 0.01
$\epsilon=0.2$ 0.9 5.8 0.0 0.2 1.0 3.8 19.7 0.1 2.6 3.9
$\epsilon=0.4$ 1.3 8.0 0.0 0.4 1.4 5.0 23.7 0.3 3.5 5.2
$\epsilon=0.6$ 1.7 10.7 0.0 0.5 1.9 6.1 27.5 0.6 4.2 6.4
$\epsilon=0.8$ 2.1 12.9 0.0 0.8 2.3 6.8 29.6 0.7 5.1 7.0
C = 0.1
$\epsilon=0.2$ 0.0 0.4 0.0 0.0 0.1 0.8 5.7 0.0 0.5 0.8
$\epsilon=0.4$ 0.2 1.9 0.0 0.1 0.3 2.3 13.4 0.1 1.4 2.4
$\epsilon=0.6$ 0.5 3.6 0.0 0.1 0.6 3.3 18.6 0.0 2.2 3.4
$\epsilon=0.8$ 0.6 4.3 0.0 0.1 0.7 3.5 20.0 0.0 2.4 3.7
Table 7: HumanART Results: DP-SGD with projection.
Privacy Parameter($\epsilon$) Head Shoulder Elbow Wrist Hip Knee Ankle Mean Mean@0.1 inkoti
Finetuning
C = 0.01
$\epsilon=0.2$ 38.3 73.9 35.6 21.7 40.4 43.6 76.9 43.6 30.8 45.8
$\epsilon=0.4$ 34.7 73.4 29.1 18.3 36.6 39.7 76.0 37.9 26.9 41.4
$\epsilon=0.6$ 40.3 75.8 38.4 22.5 42.6 46.1 78.5 47.0 33.1 47.8
$\epsilon=0.8$ 38.7 74.3 36.3 21.1 40.8 44.3 77.3 44.5 32.1 45.9
C = 0.1
$\epsilon=0.2$ 31.6 69.1 24.7 17.0 33.3 36.8 72.7 33.1 26.0 38.2
$\epsilon=0.4$ 32.9 70.4 27.0 18.0 34.6 38.1 73.2 35.2 27.2 39.5
$\epsilon=0.6$ 33.6 71.3 26.9 18.4 35.4 38.6 74.3 35.2 28.2 40.0
$\epsilon=0.8$ 33.5 69.4 28.4 18.4 35.3 38.7 72.6 36.8 27.1 40.2
C = 1.0
$\epsilon=0.2$ 1.1 6.7 0.0 0.3 1.3 4.8 21.9 0.2 3.2 5.0
$\epsilon=0.4$ 11.7 41.7 2.0 3.8 12.8 18.1 52.4 7.8 11.1 19.0
$\epsilon=0.6$ 10.7 39.4 1.9 2.9 11.7 17.1 51.2 6.8 9.7 18.0
$\epsilon=0.8$ 23.5 60.4 14.2 11.1 25.2 30.2 66.2 24.2 20.2 31.5
Finetuning from scratch
C = 0.01
$\epsilon=0.2$ 1.7 8.5 0.4 0.4 1.9 4.9 19.5 1.1 2.9 5.2
$\epsilon=0.4$ 3.5 15.2 0.3 1.1 3.9 8.4 29.0 2.6 5.8 8.7
$\epsilon=0.6$ 2.1 10.2 0.1 0.7 2.4 5.9 22.2 1.4 3.4 6.2
$\epsilon=0.8$ 1.4 7.3 0.1 0.5 1.6 4.3 18.2 0.6 2.4 4.5
C = 0.1
$\epsilon=0.2$ 0.5 2.8 0.1 0.2 0.6 2.2 10.9 0.1 1.5 2.3
$\epsilon=0.4$ 0.6 4.0 0.0 0.3 0.7 2.5 12.9 0.1 1.7 2.6
$\epsilon=0.6$ 1.2 5.7 0.5 0.4 1.5 3.4 15.3 0.6 1.7 3.7
$\epsilon=0.8$ 1.1 4.2 1.0 0.3 1.2 3.1 14.0 0.5 1.8 3.2
C = 1.0
$\epsilon=0.2$ 0.1 0.9 0.0 0.0 0.2 0.9 5.6 0.0 0.6 0.9
$\epsilon=0.4$ 0.3 2.0 0.1 0.0 0.4 1.9 11.1 0.1 1.0 2.0
$\epsilon=0.6$ 0.8 4.6 0.0 0.1 0.9 3.1 16.1 0.1 1.7 3.3
$\epsilon=0.8$ 0.2 1.3 0.1 0.0 0.3 1.3 7.5 0.1 0.9 1.4
Training from scratch
C = 0.01
$\epsilon=0.2$ 0.12 0.68 0.06 0.0 0.17 1.06 5.7 0.1 0.5 1.1
$\epsilon=0.4$ 0.0 0.09 0.0 0.0 0.0 0.09 0.71 0.0 0.10 0.09
$\epsilon=0.6$ 0.31 2.1 0.0 0.07 0.4 1.8 11.7 0.01 1.18 1.93
$\epsilon=0.8$ 0.11 0.82 0.0 0.04 0.2 1.2 7.0 0.02 1.15 1.20
C = 0.1
$\epsilon=0.2$ 0.05 0.29 0.0 0.06 0.38 2.5 0.0 0.3 0.4 1.24
$\epsilon=0.4$ 0.12 0.91 0.0 0.05 0.16 1.35 7.58 0.04 1.3 1.37
$\epsilon=0.6$ 0.19 1.27 0.0 0.04 0.23 1.6 9.06 0.03 1.02 1.6
$\epsilon=0.8$ 0.02 0.11 0.0 0.0 0.03 2.09 0.0 0.25 0.32 0.31
Table 8: HumanART Results: Feature Projection DP-SGD.
Privacy Parameter($\epsilon$) Head Shoulder Elbow Wrist Hip Knee Ankle Mean Mean@0.1 inkoti
Finetuning
C = 0.01
$\epsilon=0.2$ 32.6 71.8 26.4 17.1 34.6 38.1 75.2 35.1 26.6 39.6
$\epsilon=0.4$ 37.3 74.6 34.2 20.7 39.5 42.6 77.7 42.4 30.4 44.2
$\epsilon=0.6$ 39.3 75.9 37.6 22.2 41.7 44.5 78.8 45.4 31.4 46.2
$\epsilon=0.8$ 40.5 76.9 39.4 23.1 42.7 45.6 79.4 47.0 32.5 47.3
C = 0.1
$\epsilon=0.2$ 8.7 37.0 1.1 3.1 9.4 14.1 48.7 2.9 9.6 14.7
$\epsilon=0.4$ 13.5 47.2 2.1 4.8 14.6 18.8 55.9 6.8 12.3 19.6
$\epsilon=0.6$ 15.8 51.9 3.5 5.9 17.0 21.1 59.1 9.3 14.1 22.0
$\epsilon=0.8$ 17.1 53.5 4.6 6.6 18.4 22.4 60.6 10.8 14.9 23.3
Finetuning from scratch
C = 0.01
$\epsilon=0.2$ 4.8 23.0 0.2 0.9 5.5 10.4 38.4 2.4 6.3 10.9
$\epsilon=0.4$ 7.7 30.0 1.7 1.7 8.5 13.6 43.1 5.0 8.5 14.2
$\epsilon=0.6$ 8.8 32.8 2.3 2.0 9.7 14.8 45.2 6.3 9.3 15.4
$\epsilon=0.8$ 9.5 34.5 2.6 2.2 10.6 15.7 46.9 7.1 10.1 16.3
C = 0.1
$\epsilon=0.2$ 0.2 1.9 0.0 0.0 0.3 2.3 13.6 0.0 1.4 2.5
$\epsilon=0.4$ 0.7 4.9 0.0 0.0 0.8 3.2 18.3 0.09 1.7 3.3
$\epsilon=0.6$ 1.0 7.0 0.0 0.2 1.1 3.9 21.0 0.1 2.7 4.0
$\epsilon=0.8$ 1.3 9.3 0.0 0.2 1.5 4.4 23.1 0.1 2.8 4.6
Table 9: HumanART Results: Feature Projection DP-SGD plus projection.
Privacy Parameter($\epsilon$) Head Shoulder Elbow Wrist Hip Knee Ankle Mean Mean@0.1 inkoti
Finetuning
C = 0.01
$\epsilon=0.2$ 50.6 80.2 53.4 31.3 52.8 55.3 82.4 59.1 41.3 57.2
$\epsilon=0.4$ 51.6 80.4 55.0 33.5 53.8 56.2 82.7 60.3 42.2 58.1
$\epsilon=0.6$ 50.9 80.2 53.6 32.5 53.2 55.9 82.7 59.9 42.1 57.8
$\epsilon=0.8$ 51.6 80.2 54.7 33.6 53.7 56.2 82.5 60.2 42.3 58.1
C = 0.1
$\epsilon=0.2$ 40.4 75.5 38.9 24.4 42.3 45.4 78.0 46.0 33.0 47.1
$\epsilon=0.4$ 41.1 76.8 40.2 24.6 43.3 46.8 79.5 48.2 34.4 48.4
$\epsilon=0.6$ 42.2 75.7 42.2 25.2 44.3 47.5 78.6 49.3 34.4 49.2
$\epsilon=0.8$ 43.0 77.0 42.9 25.7 45.1 48.4 79.9 50.3 35.7 50.1
C = 1.0
$\epsilon=0.2$ 26.8 62.1 19.0 13.5 28.4 32.9 67.0 28.6 23.2 34.2
$\epsilon=0.4$ 32.4 70.5 25.4 19.0 34.2 39.0 74.7 36.6 29.9 40.3
$\epsilon=0.6$ 33.9 70.6 28.7 18.3 35.8 40.1 74.4 38.5 28.8 41.6
$\epsilon=0.8$ 35.4 72.1 30.7 20.2 37.2 40.7 75.1 38.6 29.9 42.1
Finetuning from scratch
C = 0.01
$\epsilon=0.2$ 45.3 76.9 46.4 29.5 47.3 50.4 79.7 53.0 38.0 52.1
$\epsilon=0.4$ 45.1 76.6 46.5 28.0 47.2 50.5 79.1 53.5 36.8 52.3
$\epsilon=0.6$ 43.8 75.2 44.5 28.5 45.6 49.6 78.0 52.2 38.3 51.1
$\epsilon=0.8$ 46.0 76.7 47.5 29.3 48.1 51.4 79.7 54.5 38.0 53.2
C = 0.1
$\epsilon=0.2$ 6.8 24.4 2.3 2.8 7.4 11.6 35.9 5.2 8.8 12.0
$\epsilon=0.4$ 22.0 53.7 14.5 12.0 23.3 27.7 60.3 22.1 20.6 28.7
$\epsilon=0.6$ 29.1 62.4 24.0 17.7 30.8 34.9 67.8 32.0 26.2 36.1
$\epsilon=0.8$ 33.0 68.0 27.9 21.5 34.4 38.6 72.0 36.5 30.0 39.8
Figure 4: Figures (a-e)Depiction of qualitative results on DP-SGD, Projection DP-SGD and Feature Projection DP-SGD. We specifically show results on Finetuning with $C=0.1$ at various privacy budgets. (f) Representation of Raw (Private) image compared to public feature (gaussian blurred).